OpenStudy (yanasidlinskiy):

How would I figure this out? Find a quadratic function to model the values in the table. Predict the value of y for x = 6.

OpenStudy (yanasidlinskiy):

 I've never done these before. So i really need help!

OpenStudy (lasttccasey):

Well it sounds like you need to use quadratic interpolation which is a long winded explanation. This site lays out the process so take a look at it and see if that doesn't help first. http://isezen.com/2012/01/15/quadratic-interpolation-three-point/

OpenStudy (yanasidlinskiy):

Umm. i don't really understand.

OpenStudy (lasttccasey):

Well basically you will have 3 equations with 3 unknowns. Set up goes like this: Here is the template equation:\[y = a0 + a1*x + a2*x^2\]Now the three unknowns you solve for are a0, a1, and a2. To do this you use the given data and plug it in for the three data points. These are the three equations: \[2 = a0 + a1*(1) + a2*(1)^2\]\[2 = a0 + a1*(0) + a2*(0)^2\]\[-10 = a0 + a1*(3) + a2*(3)^2\]Three equations and three unknowns can be solved by hand or using matrix algebra.

OpenStudy (lasttccasey):

Does that help? @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Oh, Ok. Yea. It starts to make a little sense now. Do you mind continuing?

OpenStudy (amistre64):

the data provides a simpler way

OpenStudy (amistre64):

since there are 2 x values with the same y value, the axis of symmetry has to be the middle between them

OpenStudy (lasttccasey):

Well no I don't mind but that's all there is to it really. Once you get a0, a1 and a2 you will use the template equations, plug in the values, and you will have f(x) = y = #x^2 + #x + #. If you want to evaluate at x = 6, f(6) = the predicted value as asked.

OpenStudy (amistre64):

since x=0 and y=2, we know the y intercept is 2

OpenStudy (lasttccasey):

Yes there's usually a 'simpler' way but this way will always work without fail.

OpenStudy (amistre64):

yeah, it will work

OpenStudy (yanasidlinskiy):

So would the equation be like..\[-2x^2+2x-2; -58\] ??

OpenStudy (amistre64):

-2 on the start looks fine, but the end should be +2, not sure about the middle yet

OpenStudy (amistre64):

yeah, middles a +2x works well

OpenStudy (yanasidlinskiy):

Ok. So are you sure that the end should be +2. Correct?

OpenStudy (amistre64):

well, when x=0, y=2 ... not -2 so yeah

OpenStudy (yanasidlinskiy):

Here's my answer choices: A. y = -2x2 + 2x -2; –58 B. y = 2x2 – 2x -2; 60 C. y = 2x2 – 2x -2; 58 D. y = -2x2 + 2x +2; –58

OpenStudy (amistre64):

only one of them fits for the point (0,2)

OpenStudy (amistre64):

unless your table data is wrong :/

OpenStudy (yanasidlinskiy):

Nope. None of it is wrong:D

OpenStudy (yanasidlinskiy):

Its the last one. Right?

OpenStudy (lasttccasey):

Good job, yes the last one is correct. Just to revisit the other method for future use. If you did interpolation, you'd find a0, a1, and a2 came out to be 2, 2, and -2 respectively. Thus the equation is y = - 2*x^2 + 2*x + 2 evaluating at 6 gives y(6) = -58

OpenStudy (yanasidlinskiy):

Thank you sooo much!! @lasttccasey and @amistre64 I appreciate your help!!:D I'm definitely writing this down in my notebook for Algebra 2!!:D

OpenStudy (amistre64):

good luck :)

OpenStudy (yanasidlinskiy):

Thanx:D